# Predicted High-Pressure Hot Superconductivity in Li2CaH16 and Li2CaH17 Phases that Resemble the Type-II Clathrate Structure

### v1

@@ 4 Conclusions
Previous theoretical calculations have predicted that “hot superconductors”, compounds that are superconducting at room temperature and above, may be stabilized under extreme pressures 41, 30, 43. Interest in these high-energy-density quantum materials stems from the fact that quantum behavior at atmospheric pressures is typically only observed at low temperatures. Herein, crystal structure prediction searches combined with density functional theory calculations predict two new hot superconductors, with critical temperatures, s, as high as 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17. These ternary hydride phases are predicted to lie on the 300 GPa convex hull (when zero-point-energy effects are included). Li2CaH17 was found to be dynamically stable down to 160 GPa in the harmonic approximation, where its was predicted to be similar to previously-synthesized H3S 3 and CaH68, 9, 10.
The structures of Li2CaH16 and Li2CaH17 can both be described as clathrate-like hydrogenic lattices where the heavier calcium cation occupies the larger H28 cages, and the smaller lithium atom is situated within H18 or H20 encapsulations. Both of these phases are related to the Type-II clathrate structure with Li2CaH17 being a rhombohedral distortion of the symmetry lattices found in the silicon variants, and Li2CaH16 being derived from the ideal Type-II clathrate structure via removal of a single hydrogen atom. We postulate that the in these ternary hydrides could be enhanced by tuning the average valence of the enclathrated electropositive metal atoms so that the Fermi level sits on the peak in the densities of states, resulting in compositions such as Li2Ca0.6M0.4H16/17, where M is a trivalent metal ion.
@@ 5 Acknowledgements
Funding for this research was provided by the NSF award DMR-2136038 and the U.S. Department of Energy, National Nuclear Security Administration, through the Capital-DOE Alliance Center under Cooperative Agreement DE-NA0003975. EZ acknowledges the U.S. Department of Energy, Office of Science, Fusion Energy Sciences funding the award entitled High Energy Density Quantum Matter under Award No. DE-SC0020340. Calculations were performed at the Center for Computational Research at SUNY Buffalo (https://hdl.handle.net/10477/79221).

### v2

@@ 4 Conclusions
Previous theoretical calculations have predicted that “hot superconductors”, compounds that are superconducting at room temperature and above, may be stabilized under extreme pressures 42, 31, 44. Interest in these materials stems from the fact that quantum behavior at atmospheric pressures is typically only observed at low temperatures. Herein, crystal structure prediction searches combined with density functional theory calculations predict two new hot superconductors, with critical temperatures, s, as high as 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17. These ternary hydride phases are predicted to lie on the 300 GPa convex hull (when zero-point-energy effects are included). Li2CaH17 was found to be dynamically stable down to 160 GPa in the harmonic approximation, where its was predicted to be similar to previously-synthesized H3S 4 and CaH69, 10, 11.
The structures of Li2CaH16 and Li2CaH17 can both be described as clathrate-like hydrogenic lattices where the heavier calcium cation occupies the larger H28 cages, and the smaller lithium cation is situated within H18 or H20 encapsulations. Both of these phases are related to the Type-II clathrate structure with Li2CaH17 being a rhombohedral distortion of the symmetry lattices found in the silicon variants, and Li2CaH16 being derived from the ideal Type-II clathrate structure via removal of a single hydrogen atom. We postulate that the in these ternary hydrides could be enhanced by tuning the average valence of the enclathrated electropositive metal atoms so that the Fermi level sits on the peak in the densities of states, resulting in compositions such as Li2Ca0.6M0.4H16/17, where M is a trivalent metal ion.
@@ 5 Acknowledgements
Funding for this research was provided by the NSF award DMR-2136038 and the U.S. Department of Energy, National Nuclear Security Administration, through the Chicago-DOE Alliance Center (CDAC) under Cooperative Agreement DE-NA0003975. E.Z. acknowledges the U.S. Department of Energy, Office of Science, Fusion Energy Sciences funding the award entitled High Energy Density Quantum Matter under Award No. DE-SC0020340. Calculations were performed at the Center for Computational Research at SUNY Buffalo (https://hdl.handle.net/10477/79221). M.R. would like to thank Dr. Francesco Belli for discussions regarding quantum fluctuations and electron-phonon coupling.
6 Supporting Information Available
The Supporting Information is available free of charge on the publication website,
http://pubs.acs.org/. It includes the space group at various pressures, distance above the convex hull, relative enthalpies, equation of states, electronic band structures and densities of states, electron localization functions, Bader charges, phonon density of states, computational parameters in EPC calculations, plots of first-principle molecular dynamic simulation results, superconducting properties of selected phases, and structural parameters. The optimized structures were deposited in as cif files in the Theoretical Crystallography Open Database (TCOD IDs: 30000120-30000153).